you have a 2Ω, a 3Ω, and a 4Ω resistor in series with a 4.5V battery. Sketch the circuit. What is the total current going through the battery?
1
Expert's answer
2019-05-20T13:55:09-0400

Answer:


The sketch of the circuit is shown on the figure 1, where

"R_1=2\Omega, R_2 = 3\Omega, R_3 = 4\Omega,"

and

"E = 4.5V."

In series circuit, current remains same in all resistances and battery. The equivalent resistance of the circuit is the next sum

"R = R_1+R_2+R_3 = 2+3+4 = 9\Omega."

The current flows through the battery is

"I=E\/R = 4.5\/9 = 0.5A."